OpenStudy (anonymous):

I am struggling with this equation 7= 10(1+e)^x/2 Can anybody help me with it?

OpenStudy (anonymous):

what about this equation?

OpenStudy (anonymous):

is it \[7=10(1+e)^{x/2}\]?

OpenStudy (anonymous):

okay, I will just solve the equation written above: first divide both sides by 10, and then take "ln" (natural logarithm) for both sides as: \[{7 \over 10}=(1+e)^{x/2} \implies \ln(0.7)={x \over 2}\ln(1+e)\] divide both sides by ln(1+e), and multiply by 2: \[{\ln(0.7) \over \ln(1+e)}={x \over 2} \implies x=2{\ln(0.7) \over \ln(1+e)}\] if this is a MC question, you can just simplify your answer according to the choices that you have.